# Measurement of Sudden Beam Loss Events Using Bunch-by-Bunch BPMs at SuperKEKB

### v1

@@ I Introduction
The SuperKEKB accelerator [1] collides 4 GeV positrons with 7 GeV electrons at extremely high luminosity, producing large numbers of B mesons, D mesons, $\tau$ leptons, etc. for the Belle II experiment [2]. In December 2024, SuperKEKB achieved a world-record instantaneous luminosity of $5.1\times 10^{34}~{}\mathrm{cm}^{-2}\mathrm{s}^{-1}$. The long-term goal is to exceed this record by more than an order of magnitude. However, a phenomenon known as Sudden Beam Loss (SBL) has emerged as a major obstacle to stable accelerator operation [3, 4].
@@
SBL refers to the abrupt loss of a large portion of the stored beam within an extremely short time—on the order of tens of microseconds—ultimately resulting in a beam abort. Since the time from beam loss detection by loss monitors to the triggering of a beam abort is typically 10–20 µs (equivalent to about 1–2 turns), the beam cannot be aborted in time before a large beam loss when an SBL event occurs. As a result, the beam strikes the Belle II detector and other accelerator components, causing severe damage. Examples include damage to the innermost silicon detectors of Belle II, collimator heads, and quenching of the superconducting final focus system located near the collision point. Addressing this issue has become an urgent priority. In order to understand the mechanism of SBL, it is essential to capture the accompanying beam instabilities in detail. We approached this challenge by focusing on bunch-by-bunch beam position oscillations measured with high-speed beam position monitors.
@@
To enable precise observation and analysis of SBL events, we have developed a new Bunch Oscillation Recorder (BOR) using AMD/Xilinx Radio Frequency System on Chip (RFSoC) technology [5]. The BOR is a beam diagnostic system that records bunch-by-bunch beam position and charge immediately before a beam abort. It enables us to observe the behavior of individual bunches as they lose charge during an SBL event. Although multiple BOR types are installed at SuperKEKB, in this paper, we refer specifically to the newly developed RFSoC-based BOR simply as “BOR.”
This paper presents our observations of SBL events using the BOR and the subsequent analysis. Section 2 describes the experimental setup used for SBL observation. Section 3 provides analyses of the location of bunch charge loss, as well as the amplitude and duration of bunch position oscillations. Section 4 discusses the correlation between bunch position oscillations and other related phenomena such as beam losses observed by the loss monitors and pressure burst events [6]. Section 5 provides concluding remarks.
II Methodology of SBL Observation
@@
Beam current $I$ [A] 1.632 1.259
Num of bunches $n_{b}$ 2346
Bunch current $I_{b}$ [mA] 0.696 0.537
Horizontal beta function at IP $\beta_{x}^{\ast}$ [mm] 60 60
Vertical beta function at IP $\beta_{y}^{\ast}$ [mm] 1.0 1.0
@@
Vertical tune $\nu_{y}$ 46.585 43.5985
Horizontal emittance $\epsilon_{x}$ [nm] 3.0 7.0
Vertical emittance $\epsilon_{y}$ [pm] 100 100
Bunch length $\sigma_{z}$ [mm] $\sim 6$ $\sim 6$
Instantaneous luminosit $L$ [$\rm{cm}^{-2}\rm{s}^{-1}$] 5.1×$\times 10^{34}$
Table 1: Main parameters of the SuperKEKB main ring. Parameters below the beam current correspond to the conditions at the time of the highest recorded luminosity on December 27, 2024.
@@ II.1 Collimator
II.2 Vacuum gauge
To monitor vacuum pressure inside the beam pipes, approximately 600 cold cathode gauges (CCGs) are installed along the main ring at intervals of about 10 m, with roughly 300 in each ring [9]. During operation, the vacuum level is maintained at $10^{-8}$ to $10^{-7}$ Pa. Each CCG is labeled with a name such as “D01_L01,” where the suffix “L” or “H” denotes whether the gauge is installed in the LER or HER, followed by a serial number indicating its position along the beamline.
II.3 Loss monitor and beam abort system
@@
II.4 SBL observation setup
Two BORs were developed and installed in the LER to observe SBL events [5]. In this paper, we focus on SBL events occurring in the LER, as previous observations have shown that SBL events occur more frequently in the LER than in the HER [14]. One BOR was installed in the Fuji straight section (bottom of Fig. 1), referred to as “Fuji-RFSoC.” The second was installed in the D05 section (right of Fig. 1), referred to as “D5-RFSoC.” These locations are marked with green triangles in Fig. 1. Between the two BORs lies the beam dump (blue star) and the D06 collimator section, which houses two horizontal collimators (D06H3 and D06H4) and two vertical collimators (D06V1 and D06V2). Upstream of Fuji-RFSoC, there is the D10 section, where wiggler magnets are installed. The D10 section contains multiple CCGs spaced about 10 m apart, named sequentially from upstream (e.g., D10_L01, D10_L02, …). Because the LER beam travels counterclockwise, the beamline proceeds from the D10 section through the Fuji-RFSoC, the beam dump, the D06 collimator section, the D5-RFSoC, and finally to the collision point. Table 2 summarizes the optics parameters at relevant locations, including CCGs in the D10 section, BOR locations, and collimators in the D06 section.
Location $s$ (m) $\beta_{x}$ (m) $\nu_{x}$ (rad/$\pi$) $\eta_{x}$ (mm) $\beta_{y}$ (m) $\nu_{y}$ (rad/$\pi$)
@@ II.4 SBL observation setup
In this paper, a beam abort is classified as an SBL event if the maximum charge loss among the bunches exceeded 5%. The Fuji-RFSoC recorded 117 SBL events between October 13 and November 28, 2024. The D5-RFSoC recorded 58 events between October 26 and November 28, 2024. During this period, SuperKEKB operated in two optics configurations. From October 13 to 23, a “Detuned Optics” setting was used, with relaxed beta functions at the collision point ($\beta_{y}^{*}=48.6$ mm, $\beta_{x}^{*}=384$ mm), and no collisions were performed. After October 24, the optics were switched to enable collisions, with the vertical beta function squeezed to $\beta_{y}^{*}=1$ mm. The collimator apertures also varied depending on the optics setting. Typical values for each setting are shown in Table 3. When the beta functions are squeezed, the collimator apertures are reduced accordingly. Since the horizontal beam size is generally larger than the vertical size, the horizontal collimators are opened wider.
D02V1
D02H1
D06V1
D06H4
Detuned optics 13 20 9 30
$\beta_{y}^{\ast}=1~{}\mathrm{mm}$ optics
2 14 4 20
Table 3: Typical collimator apertures for the two main optics settings of the SuperKEKB main ring. The collimator names correspond to those shown in Fig. 1. The values represent the distance between opposing collimator heads. In practice, these apertures are fine-tuned around the listed values during operation.
The BORs are connected to the beam abort kicker trigger (see Section II.3). Upon receiving the trigger, the BOR halts its ring buffer and stores the recorded data. This allows capture of the beam position and charge of all bunches in the main ring for 100 turns prior to the beam abort [5].
III BOR data analysis for SBL events
@@
III.1 Example of SBL observations
Figure 2 shows a representative example of an SBL event. This event occurred at 2:55:30 on October 28, 2024, with 2346 bunches circulating in the ring and an average bunch current of 0.64 mA. In this plot, each vertical division corresponds to one beam revolution (i.e., approximately 10 µs), and each dot represents a single bunch. The BOR records each bunch once per turn, so the same bunch appears repeatedly in the plot. The horizontal axis indicates the number of turns before the beam abort, with the abort occurring at the right edge of the plot. Note that in the last turn before the abort, bunches pass through the location of the Fuji-RFSoC and then enter the beam dump (see Fig. 1). Therefore, the final turn data is missing from the D5-RFSoC plot since it is located downstream of the beam dump. In Fig. 2, two bunch trains and two abort gaps can be identified within each revolution. To isolate the oscillatory component due to beam instability leading to SBL, we subtract the bunch positions recorded in the first turn (100 turns before the abort) from the position of each bunch in all turns. This removes offsets due to the nominal beam orbit. For bunch charge, we normalize the values by defining the charge in the first recorded turn (100 turns before the abort) as 100% for each bunch. This normalization allows clearer visualization of charge loss, because the bunch charges in the main ring are not perfectly uniform. All subsequent plots in this paper follow the same conventions for representing bunch position and charge. In this SBL event, oscillations in bunch positions began before any noticeable charge loss, both in the Fuji-RFSoC and D5-RFSoC data. The beam abort was ultimately triggered with a maximum bunch charge loss of approximately 40%.
@@ III.1 Example of SBL observations
III.2 Bunch charge loss
@@
The total charge loss in Section 1 is calculated by summing the orange points in Fig. 5 from 8 turns before the abort up to 1 turn before. For Section 2, the total loss is calculated by summing the blue points from 7 turns before the abort up to the final recorded turn. Figure 6 shows a histogram of the ratio $P$ calculated for 58 SBL events observed simultaneously by both BORs. In all cases, more than 50% of the charge loss occurred in Section 1. This indicates that the majority of SBL-related charge loss can be attributed to beam interaction with collimators in the D06 section. This, in turn, implies that large position oscillations or possibly an increase in beam size tend to occur before the beam enters the D06 section.
Figure 6: Distribution of the ratio $P$ (defined in Eq. 1) for 58 SBL events recorded simultaneously by the two BORs.
III.3 Duration of bunch position oscillation
@@
III.4 Amplitude of Bunch Position Oscillation
We next investigate the amplitude of bunch position oscillations during SBL events. Figure 10 illustrates the procedure used to determine the oscillation amplitude. Bunches that have already lost charge are excluded from the calculation, since partial loss due to interactions with collimators can shift the center of charge and distort the position measurement. Here, we define the amplitude as the difference between the maximum and minimum values of the bunch position recorded before the onset of charge loss. The onset of charge loss is defined as the moment when the charge of any bunch decreases by more than 5%. Using this definition, we calculate the oscillation amplitudes for all SBL events recorded by the Fuji-RFSoC and the D5-RFSoC. The resulting distributions are shown in Fig. 11. In addition to the raw amplitude (in mm), we present the “normalized amplitude,” defined as the amplitude divided by the square root of the beta function ($\sqrt{\mathrm{m}}$) at the BOR location. This normalization provides a more intrinsic measure of oscillation strength that is independent of the observation point. The normalized amplitudes are displayed on the top axes of the histograms.
@@ III.4 Amplitude of Bunch Position Oscillation
Figure 11: Distributions of bunch oscillation amplitude. Left: horizontal oscillation. Right: vertical oscillation. The top axis in each plot shows the normalized amplitude, obtained by dividing the amplitude (in mm) by the square root of the beta function ($\sqrt{\mathrm{m}}$) at the BOR location.
Focusing on the normalized amplitudes in the vertical oscillation, the results indicate that the Fuji-RFSoC tends to exhibit larger normalized amplitudes than the D5-RFSoC. This difference likely arises from our use of only pre-charge-loss data in the amplitude calculation. When strong bunch oscillations are observed at the Fuji-RFSoC, similar amplitudes would be expected at the D5-RFSoC under normal conditions. However, if beam loss occurs at the D06 collimator section immediately downstream of the Fuji-RFSoC, those bunches are no longer included in the amplitude calculation at the D5-RFSoC. As a result, the calculated amplitudes at the D5-RFSoC tend to be smaller than those at the Fuji-RFSoC. This observation suggests that strong bunch oscillations, or possibly an increase in beam size, often develop before the beam enters the D06 section and hit the collimators there—consistent with the discussion in Section III.2.
IV Analysis Using Multiple Monitoring Systems

### v2

@@ 1 Introduction
The SuperKEKB accelerator [18] collides 4 GeV positrons with 7 GeV electrons at extremely high luminosity, producing large numbers of B mesons, D mesons, $\tau$ leptons, etc. for the Belle II experiment [8]. In December 2024, SuperKEKB achieved a world-record instantaneous luminosity of $5.1\times 10^{34}~\mathrm{cm}^{-2}\mathrm{s}^{-1}$. The long-term goal is to exceed this record by more than an order of magnitude. However, a phenomenon known as Sudden Beam Loss (SBL) has emerged as a major obstacle to stable accelerator operation [6].
@@
SBL refers to the abrupt loss of a large portion of the stored beam within tens of microseconds, ultimately resulting in a beam abort. Since the response time of the abort system is typically 10–20 µs (about 1–2 turns), the beam cannot be aborted quickly enough, and significant beam loss strikes accelerator components before the abort is completed. Such uncontrolled losses have caused serious damage, including to the innermost silicon detectors of Belle II and collimator heads, as well as quenching of the superconducting final-focus magnets located near the collision point. Understanding the mechanism of SBL is therefore essential for both machine protection and future luminosity upgrades.
So far, two main diagnostics have been used to study SBL: loss monitors, which detect radiation from lost particles, and vacuum gauges, which measure pressure bursts [20] inside the beam chamber. Pressure bursts are observed in most SBL events and are thought to occur at locations suspected to be the source points of SBL [6]. While indispensable, these devices cannot capture the detailed dynamics of individual bunches leading up to an SBL event. Loss monitors integrate over many bunches and locations, and vacuum gauges have response times too slow to resolve the microsecond-scale evolution of SBL. As a result, the underlying processes—such as the onset of bunch oscillations or the sequence of charge depletion—remain not well understood.
@@
To overcome these limitations, we have developed a new Bunch Oscillation Recorder (BOR) using AMD/Xilinx Radio Frequency System-on-Chip (RFSoC) technology [13]. The BOR is a dedicated diagnostic that records bunch-by-bunch beam position and charge for all stored bunches during the final $\sim$100 turns before a beam abort. This enables direct observation of individual bunch dynamics during SBL events, which has not been possible with conventional diagnostics. Although multiple BOR types are installed at SuperKEKB, in this paper we refer specifically to the newly developed RFSoC-based BOR, hereafter referred to simply as “BOR.”
The main purpose of this study is to demonstrate the diagnostic capability of the BOR — to verify whether it can serve as a new “eye” for observing fast beam instabilities and to clarify what new information it can reveal that has been inaccessible with conventional BPMs, loss monitors, or vacuum gauges. Through direct observations of SBL events, we show that the BOR provides complementary and previously unattainable information compared with these existing diagnostics. With the BOR, we can identify when and where bunch oscillations start, quantify their amplitudes and durations, and determine in which section of the ring the charge loss occurs. By correlating these results with simultaneous measurements from loss monitors and vacuum gauges, we obtain new insights into the causal chain linking pressure bursts, beam instabilities, and sudden beam losses. This paper presents the first systematic observations and analyses of SBL events using the BOR at SuperKEKB, establishing its role as a powerful and unique diagnostic tool that deepens our understanding of the temporal evolution of SBL phenomena and provides valuable perspectives for future beam instability studies.
At the Large Hadron Collider (LHC), unidentified falling objects (UFOs)—micrometer-sized dust particles interacting with the circulating beam—had been reported to cause rapid beam losses [3, 9, 1]. Although the beam particle species accelerated at SuperKEKB and the LHC are different, it is important to examine whether the underlying mechanism is similar to that of dust-induced losses or fundamentally different. No previous study has analyzed such fast beam-loss phenomena using bunch-by-bunch position data. This work therefore not only sheds light on the mechanisms behind SBL at SuperKEKB but also provides insights of potential relevance to other accelerator facilities.
This paper is organized as follows. Section 2 describes the experimental setup for SBL observation. Section 3 presents analyses of SBL events recorded by the BOR, including charge loss distributions, oscillation onset times, and oscillation amplitudes. Section 4 combines BOR results with loss monitors and vacuum gauges, highlighting correlations with pressure bursts and bunch oscillations. Finally, Section 5 summarizes the key findings and discusses prospects for further studies.
2 Methodology of SBL Observation
@@
Beam current II [A] 1.632 1.259
Number of bunches $n_{b}$ 2346
Charge per 1 mA bunch [nC] 10
Horizontal beta function at IP $\beta_{x}^{\ast}$ [$m$] 60 60
Vertical beta function at IP $\beta_{y}^{\ast}$ [$m$] 1.0 1.0
@@
Vertical tune $\nu_{y}$ 46.589 43.599
Vertical beam size at IP $\sigma_{y}^{\ast}$ [nm] $\sim 300$
Bunch length $\sigma_{z}$ [$m$] $\sim 6$
[…]
Table 1: Main parameters of the SuperKEKB main ring. Parameters below the beam current correspond to the conditions at the time of the highest recorded luminosity on December 27, 2024 [19].
@@ 2.1 Collimator
2.2 Vacuum gauge
To monitor vacuum pressure inside the beam chambers, approximately 600 cold cathode gauges (CCGs) are installed along the main ring spaced approximately 10 m apart, with roughly 300 in each ring [21]. During operation, the vacuum level is maintained at $10^{-8}$ to $10^{-7}$ Pa. Each CCG is labeled with a name such as “D01_L01,” where the suffix “L” or “H” denotes whether the gauge is installed in the LER or HER, followed by a serial number indicating its position along the beamline.
2.3 Loss monitor and beam abort system
@@
2.4 SBL observation setup
Two BORs were developed and installed in the LER to observe SBL events [13]. In this paper, we focus on SBL events occurring in the LER, as previous observations have shown that SBL events occur more frequently in the LER than in the HER [23]. One BOR was installed in the Fuji straight section (bottom of Fig. 1), referred to as “Fuji-RFSoC.” The second was installed in the D05 section (right of Fig. 1), referred to as “D5-RFSoC.” These locations are marked with green triangles in Fig. 1. It should be noted that the number of BORs was limited to two during this study. This limitation was primarily due to budgetary constraints and operational restrictions, as measurements were permitted only at BPM ports that did not interfere with routine machine operation. Therefore, the purpose of this work is not to provide a statistical survey of SBLs over the entire ring, but rather to demonstrate how much new and detailed information can be obtained even from a small number of observation points.
Between the two BORs lies the beam dump (blue star) and the D06 collimator section, which houses two horizontal collimators (D06H3 and D06H4) and two vertical collimators (D06V1 and D06V2). Upstream of Fuji-RFSoC, there is the D10 section, where wiggler magnets are installed. The D10 section contains multiple CCGs spaced about 10 m apart, named sequentially from upstream (e.g., D10_L01, D10_L02, …). Because the LER beam travels counterclockwise, the beamline proceeds from the D10 section through the Fuji-RFSoC, the beam dump, the D06 collimator section, the D5-RFSoC, and finally to the collision point. Table 2 summarizes the optics parameters at relevant locations, including CCGs in the D10 section, BOR locations, and collimators in the D06 section [15].
[…]
@@ 2.4 SBL observation setup
In this paper, a beam abort is classified as an SBL event if the maximum charge loss among the bunches exceeds 5%. The Fuji-RFSoC recorded 117 SBL events between October 13 and November 28, 2024. The D5-RFSoC recorded 58 events between October 26 and November 28, 2024. During this period, SuperKEKB operated in two optics configurations. From October 13 to 23, a “Detuned Optics” setting was used, with relaxed beta functions at the collision point ($\beta_{y}^{*}=48.6$ $m$, $\beta_{x}^{*}=384$ mm), and no collisions were performed. After October 24, the optics were switched to enable collisions, with the vertical beta function squeezed to $\beta_{y}^{*}=1$ mm. The collimator apertures also varied depending on the optics setting. Typical values for each setting are shown in Table 3. When the beta functions are squeezed, the collimator apertures are reduced accordingly. Since the horizontal beam size is generally larger than the vertical size, the horizontal collimators are opened wider.
D02V1 D02H1 D06V1 D06H4
Detuned optics ($m$) 13 20 9 30
in beam sigmas 450 80 130 110
$\beta_{y}^{\ast}=1~\mathrm{mm}$ optics ($m$) 2 14 4 20
in beam sigmas 70 50 60 70
[…]
The BORs are connected to the beam abort kicker trigger (see Section 2.3). Upon receiving the trigger, the BOR halts its ring buffer and stores the recorded data. This allows capture of the beam position and charge of all bunches in the main ring for 100 turns prior to the beam abort [13].
3 BOR data analysis for SBL events
@@
3.1 Example of SBL observations
Figure 2 shows a representative example of an SBL event. This event occurred at 2:55:30 on October 28, 2024, with 2346 bunches circulating in the ring and an average bunch current of 0.64 mA. In this plot, each vertical division corresponds to one beam revolution (i.e., approximately 10 µs), and each dot represents a single bunch. The BOR records each bunch once per turn, so the same bunch appears repeatedly in the plot. The horizontal axis indicates the number of turns before the beam abort, with the abort occurring at the right edge of the plot. Note that in the last turn before the abort, bunches pass through the location of the Fuji-RFSoC and then enter the beam dump (see Fig. 1). Therefore, the final turn data is missing from the D5-RFSoC plot since it is located downstream of the beam dump. In Fig. 2, two bunch trains and two abort gaps can be identified within each revolution. To isolate the oscillatory component due to beam instability leading to SBL, we subtract the bunch positions recorded in the first turn (100 turns before the abort) from the position of each bunch in all turns. This removes offsets due to the nominal beam orbit. For bunch charge, we normalize the values by defining the charge in the first recorded turn (100 turns before the abort) as 100% for each bunch. This normalization allows clearer visualization of charge loss because the bunch charges in the main ring are not perfectly uniform. All subsequent plots in this paper follow the same conventions for representing bunch position and charge. In this SBL event, oscillations in bunch positions began before any noticeable charge loss, both in the Fuji-RFSoC and D5-RFSoC data. The beam abort was ultimately triggered with a maximum bunch charge loss of approximately 40%. As shown in Table 2, the horizontal beta function at the D5-RFSoC installation location is relatively small, so the horizontal position oscillations of D5-RFSoC appears to be small in this plot. Some oscillating bunches do not exhibit charge loss. Charge loss occurs when a portion of the beam hits either the collimator and the beam chamber, and therefore depends not only on the position oscillation but also on the beam size, oscillation amplitude, and phase. Consequently, it is trivial that some oscillating bunches do not experience charge loss.
@@ 3.1 Example of SBL observations
3.2 Bunch charge loss
@@
The total charge loss in Section 1 is calculated by summing the orange points in Fig. 5 from 8 turns before the abort up to 1 turn before. For Section 2, the total loss is calculated by summing the blue points from 7 turns before the abort up to the final recorded turn. Figure 6 shows a histogram of the ratio $P$ calculated for 58 SBL events observed simultaneously by both BORs. In all cases, more than 50% of the charge loss occurred in Section 1. This indicates that the majority of SBL-related charge loss can be attributed to beam interaction with collimators in the D06 section. Among the many collimators installed in the ring, the analysis indicates that not those located in the D02 or D05 sections but specifically the collimator in the D06 section plays a critical role during SBL events. This, in turn, implies that large position oscillations or possibly an increase in beam size tend to occur before the beam enters the D06 section.
Figure 6: Distribution of the ratio $P$ (defined in Eq. 3.1) for 58 SBL events recorded simultaneously by the two BORs.
3.3 Duration of bunch position oscillation
@@
The duration of these oscillations is often longer than the interval between the onset of beam loss and the beam abort. If the onset of bunch position oscillations can be detected in real time and used as a trigger for the abort system, it may be possible to abort the beam before a large loss occurs, thereby protecting the accelerator from radiation damage. We are currently implementing and testing additional BOR functions for oscillation detection based on this moving average method and for trigger generation [14]. In the future, the BOR may also serve as part of the accelerator protection system.
@@
3.4 Amplitude of Bunch Position Oscillation
We next investigate the amplitude of bunch position oscillations during SBL events. Figure 10 illustrates the procedure used to determine the oscillation amplitude. Bunches that have already lost charge are excluded from the calculation, since partial loss due to interactions with collimators can shift the center of charge and distort the position measurement. Here, we define the amplitude as the difference between the maximum and minimum values of the bunch position recorded before the onset of charge loss. The onset of charge loss is defined as the moment when the charge of any bunch decreases by more than 5%. Using this definition, we calculate the oscillation amplitudes for all SBL events recorded by the Fuji-RFSoC and the D5-RFSoC. The resulting distributions are shown in Fig. 11. In addition to the raw amplitude (in mm), we present the “normalized amplitude,” defined as the amplitude divided by the square root of the beta function ($\sqrt{\beta_{x,y}(\mathrm{m})}$) at the BOR location. This normalization provides a more intrinsic measure of oscillation strength that is independent of the observation point. The normalized amplitudes are displayed on the top axes of the histograms.
@@ 3.4 Amplitude of Bunch Position Oscillation
Figure 11: Distributions of bunch oscillation amplitude. Left: horizontal oscillation. Right: vertical oscillation. The top axis in each plot shows the normalized amplitude, obtained by dividing the amplitude (in $m$) by the square root of the beta function ($\sqrt{\mathrm{m}}$) at the BOR location.
Focusing on the normalized amplitudes in the vertical oscillation, the results indicate that the Fuji-RFSoC tends to exhibit larger normalized amplitudes than the D5-RFSoC. This difference likely arises from our use of only pre-charge-loss data in the amplitude calculation. When strong bunch oscillations are observed at the Fuji-RFSoC, similar amplitudes would be expected at the D5-RFSoC under normal conditions. However, if beam loss occurs at the D06 collimator section immediately downstream of the Fuji-RFSoC, those bunches are no longer included in the amplitude calculation at the D5-RFSoC. As a result, the calculated amplitudes at the D5-RFSoC tend to be smaller than those at the Fuji-RFSoC. This observation suggests that strong bunch oscillations, or possibly an increase in beam size, often develop before the beam enters the D06 section and hit the collimators there—consistent with the discussion in Section 3.2.
4 Analysis Using Multiple Monitoring Systems